# Full-Duplex-Bench: A Benchmark to Evaluate Full-duplex Spoken Dialogue Models on Turn-taking Capabilities

### v1

@@ 3.4 Data Curation
We explain how to curate the appropriate data for each evaluation dimension as follows:
  • Candor (Pause Handling, Smooth Turn-taking): We utilize Candor Reece et al. (2023), an 850-hour open-ended spontaneous speech conversation dataset with two-channel recording, as our data source. To select suitable data for our benchmark, for pause handling: we remove instances where the pause break within one turn exceeds 0.6 seconds but is shorter than 1.5 seconds. For smooth turn-taking, we calculate the gap between each pair of consecutive turns and retain only those where all gaps fall below 0.4 seconds. Since Candor does not offer an official train–test split, we randomly sample the dataset to 897 and 539 samples on pause handling and smooth turn-taking, respectively.
@@
  • ICC (Backchannel): Umair et al. (2024) use the In Conversation Corpus (ICC) to collect Transition Relevance Places (TRPs), which are points in a speaker’s utterance that signal appropriate moments for the listener to respond. This dataset consists of 28.33 minutes of high-quality informal American English dialogues, capturing responses from 118 native speakers with no prior expertise in turn-taking research. Participants were instructed to provide brief backchannel responses (e.g., “hmm,” “yes”) whenever they found it appropriate. Their responses were recorded on individual audio channels, synchronized with the stimulus audio. On average, 59 participants responded to each stimulus turn, sometimes multiple times if they perceived multiple Transition Relevance Places, leading to an average of 159 responses per turn. This allows for estimating both the likelihood of perceiving a TRP at a given moment and the distribution of these perceived response locations. By segmenting the audio into 200-ms time windows and normalizing the backchannel counts, we generate a ground truth backchannel distribution $Q$ for computing JSD for each stimulus.
@@
  • Synthetic Data (User Interruption, Pause Handling): To address the scarcity of user interruption in existing public datasets, we generated text dialogues containing both contextual and interrupting turns using GPT-4o (OpenAI, 2023). Prompts for generating these user interruption data are provided in Appendix Table 4. Speech synthesis was performed using ChatTTS [^2] https://github.com/2noise/ChatTTS is utilized to synthesize speech from text with 10 different speaker voices randomly assigned to each sample. Leveraging the pause break tag ([break]) supported by ChatTTS, we also generated synthetic pause handling data to evaluate the model’s robustness to pauses. The prompt for generating pause handling data is also found in Appendix Table 3. This process resulted in 200 samples for user interruption and 137 samples for pause handling.
@@
Metric TOR ($\downarrow$) TOR ($\downarrow$) TOR ($\downarrow$) Freq ($\uparrow$) JSD ($\downarrow$) TOR ($\uparrow$) Latency ($\downarrow$) TOR ($\uparrow$) GPT-4 ($\uparrow$) Latency ($\downarrow$)
dGSLM 0.949 0.953 0.782 0.013 0.950 0.989 0.572 0.895 0.201 3.972
Moshi 1.000 0.989 1.000 0.005 0.977 0.996 0.112 1.000 0.765 0.037
Freeze-Omni 0.672 0.287 0.782 0.002 0.984 0.369 1.168 0.795 3.371 1.200
Table 2: Models comparison. We evaluate several models across different conversational dimensions, where Latency is presented in seconds.
4 Models Under Evaluation
To fully assess full-duplex behavior, we review these models and compare their key features in Table 1. Some systems have detailed designs and are classified as Transparent Models (e.g., SyncLLM Veluri et al. (2024), Moshi Défossez et al. (2024), dGSLM Nguyen et al. (2023), VITA Fu et al. (2024), Freeze-Omni Wang et al. (2024c), SALMONN-omni Yu et al. (2024), Wang et al. (2024a), Parrot Wang et al. (2025), OmniFlatten Zhang et al. (2024a), MinMo Chen et al. (2025), and RTTL-DG Mai and Carson-Berndsen (2025)). In contrast, Closed-Source Commercial Models (e.g., GPT-4o Voice Mode [^3] https://openai.com/index/hello-gpt-4o/, Gemini Live Voice Chat [^4] https://blog.google/products/gemini/made-by-google-gemini-ai-updates/, and Doubao [^5] https://team.doubao.com/en/special/realtime_voice) are available only via apps or demos. For closed-source systems that claim full-duplex capabilities, we treat them as such despite limited design details.
@@ 4 Models Under Evaluation
Since many models do not release complete speech-to-speech checkpoints (see Table 1, column “S2S Release”), our evaluation focuses on open-source, full-duplex systems. By open-sourcing our data and metrics, we encourage the broader community, including those with closed-source models, to evaluate their systems using our benchmark. Currently, the models evaluated in this work are those with a publicly released speech-to-speech inference pipeline:
  • dGSLM Nguyen et al. (2023): dGSLM is a textless model designed to generate natural spoken dialogues directly from raw audio. Its three-stage architecture consists of: (1) a HuBERT+k-means encoder Hsu et al. (2021) that converts speech into discrete units; (2) a dual-tower Transformer-based dialogue language model with cross-attention for turn synchronization; and (3) a HiFi-GAN decoder that reconstructs speech waveforms. Trained on 2,000 hours of two-channel telephone conversations Cieri et al. (2004), dGSLM captures both verbal and non-verbal communication cues, such as laughter and backchanneling. While the original model lacks online interaction capabilities (as both two channels are predicted from context), we extend it to interact with live user speech, enabling to simulate the online user interaction. Details of this extension are provided in Appendix A.
  • Moshi  Défossez et al. (2024): Moshi treats spoken dialogue as a speech-to-speech generation task by combining a 7B text-based LLM (Helium) with a neural audio codec (Mimi) that uses residual vector quantization for speech processing. Its hierarchical design leverages an “Inner Monologue” to predict time-aligned text tokens before generating audio tokens, enhancing linguistic quality. Unlike dGSLM, Moshi processes real user speech in real time through a multi-stream design that supports overlaps and interruptions. While it serves as a general-purpose dialogue system, its complex inference process requires task-specific adaptations, which are detailed in Appendix B.
    Figure 3: Ablations on backchanneling. We vary the VAD Silero (2024) threshold for model response segmentation and evaluate backchanneling cues. The VAD threshold ranged from 0.1 to 0.6. For JSD, lower values indicate more natural responses, while lower TOR values are desirable. Higher frequency values indicate more backchannels. dGSLM consistently delivered more backchannel capabilities compared to other models.
  • Freeze-Omni (Wang et al., 2024c): Freeze-Omni is a cascaded full-duplex speech-to-speech dialogue model that achieves real-time interaction using a chunk-wise state prediction mechanism combined with a streaming speech architecture, while keeping the LLM frozen. The system utilizes a voice activity detector (VAD) to identify the start of speech and processes it in chunks using a streaming encoder. This encoder maps acoustic features into the LLM text space. A supplementary classification layer predicts the conversational state of each chunk, enabling real-time detection of whether the user is continuing the dialogue or initiating an interruption. To manage interruptions effectively, the system incorporates multiple modules functioning as monitors and speakers, facilitating full-duplex interaction. Local inference is performed using the official release server.
5 Results
@@
Table 2 presents the results on four evaluation dimensions.
Avoiding Interruptions During Speaker Pauses
@@
All three SDMs show a high Takeover Rate (TOR) when handling pauses, suggesting they often interrupt speakers during natural breaks. The end-to-end models (dGSLM and Moshi) tend to interrupt more frequently in both real and synthetic datasets. In contrast, Freeze-Omni, which includes a dedicated module for predicting speaking and listening states, produces a lower interruption rate. This indicates that a state-control component helps manage pauses more effectively, and it may benefit end-to-end systems if integrated.
Backchanneling Dynamics
We examined three models (dGSLM, Moshi, and Freeze-Omni) on the ICC corpus using TOR, backchannel frequency (Freq), and Jensen–Shannon Divergence (JSD). Similar to the pause-handling results, Moshi often takes over the turn with whisper-like responses, resulting in a high TOR. dGSLM and Freeze-Omni have lower TOR values, with dGSLM achieving the highest Freq and lowest JSD, indicating stronger backchannel behavior and a more natural conversational flow. Freeze-Omni remains largely silent, providing few backchannels.
We also tested different thresholds for the VAD Silero (2024), ranging from 0.1 to 0.6 (with a default of 0.5), to see how speech-segmentation granularity affects model performance (Table 2 and Figure 3). dGSLM shows the greatest variation, especially in Freq and JSD, while Moshi continues to speak through most turns (TOR = 1.0). Freeze-Omni is relatively stable but produces fewer backchannels. This confirms that despite segment granularity affects model performance, each model’s core tendencies remain: dGSLM backchannels actively, Moshi whispers persistently, and Freeze-Omni stays mostly silent.
Latency of Turn-Taking
We measured TOR and average response times on the Candor dataset. Both dGSLM and Moshi produce high TOR values and respond quickly. Freeze-Omni shows a higher latency due to its cascaded design, which first generates text and then uses text-to-speech. Its lower TOR may reflect missed opportunities to take over the turn because it fails to detect the end of a speaker’s segment.
Managing User Interruptions
@@
Freeze-Omni achieves the best performance in user interruption handling, with significantly higher contextual relevance scores and acceptable latency. This is likely due to its robust large language model and multi-module architecture, which help manage user-initiated interruptions. By contrast, the end-to-end SDMs produce less coherent responses. Moshi responds quickly but yields a lower GPT-4 score, while dGSLM scores poorly overall, with high latency and reduced content quality. These results suggest that fully end-to-end systems face additional challenges in maintaining semantic coherence when users interrupt.
6 Conclusion
In this paper, we introduced Full-Duplex-Bench, a new benchmark designed to evaluate key aspects of full-duplex spoken dialogue models. By focusing on critical dimensions (pause handling, backchanneling, smooth turn-taking, and user interruption management), our framework fills an important gap left by existing benchmarks that largely target half-duplex systems or rely on coarse corpus-level metrics.
Our evaluation protocol employs automatic metrics to enable systematic and reproducible assessments of real-time interaction. Experiments on open-source full-duplex models reveal both promising capabilities and clear areas for improvement, particularly in achieving natural turn transitions and appropriate response timing.
We believe that Full-Duplex-Bench will serve as a valuable resource for the research community. By open-sourcing our data and evaluation metrics, we hope to foster collaboration and drive the development of more interactive and human-like spoken dialogue systems.
Limitations and Future Works
While our benchmark marks an important step in evaluating full-duplex SLMs, it also reveals several limitations that we plan to address in future work:
  • English-Only Corpora: Currently, our benchmark data is in English. However, turn-taking behavior may vary across languages and cultures (Stivers et al., 2009), and backchanneling signal words and frequency might differ between languages (White, 1989).
  • Closed-Sourced Commercial Model: We did not evaluate the commercial models, as it only allows interaction via a microphone and does not support local inference. Reproducing results is challenging due to variations in recording environments. In the future, we plan to explore better ways to interact with such models while ensuring reproducibility and efficiency.
  • More Accurate Backchannel Detection: We currently detect backchannels based on short duration and minimal word usage. However, incorporating a more advanced detection model that integrates temporal, lexical, and prosodic signals could improve accuracy.
  • Non-lexical Expression: Spontaneous spoken dialogue includes non-lexical expressions such as laughter and sighs. While our approach transcribes speech into text using ASR, our current metrics and evaluation dimensions do not account for these non-lexical elements. Future work should incorporate such expressions to enhance the interactivity and naturalness of full-duplex models.
Appendix A dGSLM Implementation
Denote user speech $U=[U_{1},U_{2},...,U_{T}]$ as the user stream and $M=[M_{1},M_{2},...,M_{T}]$ as the model stream, where $U_{i}$, $M_{i}$ are speech tokens after tokenization (Hsu et al., 2021), and $i=1,2,...,T$. In vanilla dGSLM (Nguyen et al., 2023), the model takes a two-channel prompt $([U_{1},U_{2},...,U_{p}],[M_{1},M_{2},...,M_{p}])$ and samples full sequences $([U_{1},U_{2},...,U_{p},U_{p+1},...,U_{T}],$ $[M_{1},M_{2},...,M_{p},M_{p+1},...,M_{T}])$ in an cross modality autoregressive manner. Specifically, $U_{i},M_{i}=f_{\theta}(U_{<i},M_{<i})$. Note that under this setting, it is not straightforward to achieve online human-model interaction, as user speech $U$ is generated and then fed as input. Based on this context, we extend dGSLM to allow flexible speech interaction, and we detail each variant below. The overall paradigm is visualized in Figure 4.
dGSLM Pause Handling/Backchanneling
We apply user speech as teacher forcing and only generate model speech. That is, $M_{i}=f_{\theta}(U^{T}_{<i},M_{<i})$, where $U^{T}$ denotes ground truth user speech. Different from vanilla dGSLM, all user speech serves as input and all model speech is output. We then analyze model output speech, which is still time synchronous with user speech, to evaluate pause handling and backchanneling performance. We also detail our teacher forcing algorithm in Sec. 21. $S_{U,t},S_{M,t}$ represent scores for user and model streams at time $t$. [BOS], [EOS] represent beginning and end of sequence tokens
Algorithm 1 Teacher Forcing Generation with dGSLM
1:User speech tokens $U^{T}$, beam size $k$, max length $T$
2:Generated model speech tokens $M$
3:Initialize $M$ with padding tokens
4:Initialize scores matrix $S$ with zeros
[…]
6:for $t=1$ to $T$ do
7:     $\textit{lprobs}_{U},\textit{lprobs}_{M}\leftarrow f_{\theta}(U^{T}_{<t},M_{<t})$triangleright Get logits for both streams
8:     if $t<|U^{T}|$ then
9:         $U_{t}\leftarrow U^{T}_{t}$triangleright Teacher forcing user stream
10:         $S_{U,t}\leftarrow 0$triangleright Zero score for teacher forced tokens
11:     else
[…]
13:     end if
14:     $\textit{cand}_{M}\leftarrow\text{BeamSearch}(\textit{lprobs}_{M},k)$triangleright Search for model tokens
15:     $M_{t}\leftarrow\text{TopBeam}(\textit{cand}_{M})$triangleright Select top beam candidate
16:     $S_{M,t}\leftarrow\text{Score}(\textit{cand}_{M})$triangleright Update model scores
[…]
18:         break
19:     end if
20:end for
21:return $(M,S)$triangleright Return generated sequence and scores
dGSLM Smooth Turn Taking
Evaluation of turn-taking is implemented as a simple spoken QA task where we process user speech and await model responses. To integrate such spoken QA capability in dGSLM, we first define prefix tokens $([U_{1},U_{2},...,U_{p}],[M_{1},M_{2},...,M_{p}])$, where the user branch contains user speech while the model branch contains pure silence. We then process a new user prompt as input $[U^{T}_{p+1},...,U^{T}_{T}]$ consisting of pure silence, and apply the same teacher forcing algorithm (Sec. 21) to generate the model response: $M_{i}=f_{\theta}(U_{1:p},M_{1:p},U_{<i}^{T})$.
dGSLM User Interruption
We implement user interruption using two dGSLM modules. As shown in Figure 4, in the first stage, we apply vanilla dGSLM inference where, given a user context wav and a silent model wav, we prompt both user and model responses, keeping only the model response while setting the user response to silence. When integrating user-interrupting speech into the current conversational stream, we first adopt offline VAD Silero (2024) to detect model responses. We select the first significant segment (duration > 1s) and obtain its boundary $[T_{\text{start}},T_{\text{end}}]$. We inject the interrupting voice at time $T_{I}=(T_{\text{end}}-T_{\text{start}})/2$. Then we concatenate all historical information and form a two-channel speech prompt. Let the user context token length be $p$ and the user interrupting token length be $q$. Our user stream will be $[U_{1},...,U_{p}]$, $[U_{p+1},...,U_{T_{m}}]$, $[U_{T_{m}},...,U_{T_{m}+q},...,U_{T}]$, corresponding to context speech, padded silence, user interruption speech, and padded silence if necessary. Our model stream will be $[M_{1},...,M_{T_{\text{start}}}]$, $[M_{T_{\text{start}}+1},...,M_{T_{\text{end}}}]$, $[M_{T_{\text{end}}+1},...,M_{T}]$, corresponding to silent speech, model’s first significant response, and remaining speech (if there is more than one significant response or backchannels), with padded silence if necessary. The second dGSLM module takes the concatenated prompts $[U_{1},...,U_{T}]$ and $[M_{1},...,M_{T}]$ as input and generates a subsequent response sequence $[M_{T+1},...,M_{T^{\prime}}]$, where we retain only the model’s generated response for evaluation purposes
Figure 5: Moshi’s inference pipeline for different conversational scenarios. (a) Vanilla time-asynchronous response generation with prefilling and silent response generation, which is also the pipeline that the smooth turn taking follows. (b) Time-synchronous generation for backchanneling and pause handling, and (c) User interruption handling through controlled user speech timing, where silence and user speech are strategically input to generate appropriate model responses. All scenarios utilize a single model with 80ms chunk-wise processing.
Appendix B Moshi Implementation
Given user speech $U=[U_{1},U_{2},...,U_{T}]$, Moshi Défossez et al. (2024) predicts the model stream $M=[M_{1},M_{2},...,M_{T}]$ in a streaming manner (chunk-wise, with each chunk spanning 80ms), enabling conversational dynamics such as backchannels. Like dGSLM, Moshi can also predict the user stream, providing flexibility for general-purpose spoken dialogue systems. We detail Moshi’s adaptation to different inference tasks in Figure 5.
Vanilla Design
The fundamental task in spoken dialogue is time-asynchronous speech-to-speech response. To achieve this, we input user speech $U$ into the system. Although Moshi will also autoregressively generate time-synchronous responses in real-time, this part is ignored during Prefilling. After prefilling, silence speech is used as input, and Moshi autoregressively generates its response. In this case, there is no speech overlap. Smooth Turn Taking follows this vanilla pipeline.
Moshi Backchanneling and Pause Handling
Time-synchronous model response generation is required to explore backchanneling and pause-handling capabilities. This process is straightforward: we input speech, and the model autoregressively generates the synchronous response.
Moshi User Interrupting
User interruption modeling requires fine-grained control of user speech timing. To achieve this, we first perform prefilling (inputting user speech while ignoring model output). Unlike vanilla or smooth turn-taking designs, instead of inputting the entire silence speech user stream to generate a response, we predetermine a generation length. For the first half of the generation, we input silence; for the second half, we begin inputting the entire user interrupt speech until reaching the maximum generation length. If any user interrupts speech remains unprocessed, we feed it into the model to continue generating model speech. Finally, after fully processing the user’s interrupting speech, we input silence as a prompt to generate the model response. In contrast to dGSLM’s design (Sec. A), this process uses only one model.
Appendix C Additional Backchanneling Ablation
We also reported the standard deviation of each metric across three models. The results are displayed in Table 6. We also visualize these statistics in Figure 6.
Figure 6: VAD ablation study. We vary the VAD Silero (2024) threshold for the backchanneling behaviors of several models.
Model TOR Freq JSD
dGSLM 0.782 ± 0.413 0.013 ± 0.028 0.950 ± 0.028
Moshi 1.000 ± 0.000 0.005 ± 0.015 0.977 ± 0.067
Freeze-Omni 0.782 ± 0.413 0.002 ± 0.007 0.984 ± 0.058
Table 6: Performance comparison of models on backchanneling metrics

### v2

@@ III-C Data Curation
The overall evaluation set of Full-Duplex-Bench is shown in the Table II. We explain how we collect the data for each evaluation dimension as follows:
Candor (Pause Handling, Smooth Turn-taking): We utilize Candor [43], an 850-hour dataset of open-ended, spontaneous speech conversations with two-channel recordings, as our primary data source. To construct suitable data samples for our benchmark:
For the pause handling task, we first apply voice activity detection to identify segments with no overlapping speech, ensuring that only one speaker is active and dominant. We then select turns that contain an internal pause lasting between 0.4 and 1 second. Additionally, we filter out turns that are too short (under 5 seconds) and exclude cases where the speech before or after the pause consists of backchannel responses.
For the smooth turn-taking task, we compute the gap between each pair of consecutive turns and retain only those where all gaps are less than 0.4 seconds. We ensure that there is no overlapping speech, each turn is longer than 4 seconds, and that the speech is not backchanneling. To allow the evaluated model sufficient time to respond, we append 5 seconds of silence at the end of the input stream. After automatic filtering, we manually review the audio samples to ensure quality and consistency.
@@
ICC (Backchannel): Umair et al. [44] use the In Conversation Corpus (ICC) to collect Transition Relevance Places (TRPs), which are points in a speaker’s utterance that signal appropriate moments for the listener to respond. This dataset consists of 28.33 minutes of high-quality informal American English dialogues, capturing responses from 118 native speakers with no prior expertise in turn-taking research. Participants were instructed to provide brief backchannel responses (e.g., “hmm,” “yes”) whenever they found it appropriate. Their responses were recorded on individual audio channels, synchronized with the stimulus audio. On average, 59 participants responded to each stimulus turn. This allows for estimating both the likelihood of perceiving a TRP at a given moment and the distribution of these perceived response locations. By segmenting the audio into 200-ms time windows and normalizing the backchannel counts, we generate a ground truth backchannel distribution $Q$ for computing JSD for each stimulus.
@@
Synthetic Data (User Interruption, Pause Handling): To address the scarcity of user interruptions in existing public datasets, we generated synthetic dialogues using GPT-4o [45], which include both contextual turns and interruption turns. Text-to-Speech synthesis was performed using ChatTTS [^6] https://github.com/2noise/ChatTTS, with 10 different speaker voices randomly assigned across samples to enhance diversity. For each input stream fed to the evaluated model, the interrupting speech is played 7 seconds after the preceding utterance. Additionally, we append 15 seconds of silence after the interruption to allow the model time to respond.
For the synthetic pause handling task, we leverage the [uv_break] tag supported by ChatTTS to insert controlled pauses into the speech, enabling us to evaluate each model’s robustness to intra-turn pauses. In total, we collect 200 samples for user interruption and 137 samples for pause handling.
TABLE III: Models comparison. We evaluate several models across different conversational dimensions, where Latency is presented in seconds.
@@
Metric TOR ($\downarrow$) TOR ($\downarrow$) TOR ($\downarrow$) Freq ($\uparrow$) JSD ($\downarrow$) TOR ($\uparrow$) Latency ($\downarrow$) TOR ($\uparrow$) GPT-4o($\uparrow$) Latency ($\downarrow$)
dGSLM 0.934 0.935 0.691 0.015 0.934 0.975 0.352 0.917 0.201 2.531
Moshi 0.985 0.980 1.000 0.001 0.957 0.941 0.265 1.000 0.765 0.257
Freeze-Omni 0.642 0.481 0.636 0.001 0.997 0.336 0.953 0.867 3.615 1.409
Gemini Live 0.255 0.310 0.091 0.012 0.896 0.655 1.301 0.891 3.376 1.183
IV Models Under Evaluation
@@
Since many models do not release complete speech-to-speech checkpoints (see Table I, column “S2S Release”) [^7] By open-sourcing our data and metrics, we encourage the broader community, including those with closed-source models, to evaluate their systems using our benchmark, currently we evaluate models with a publicly available speech-to-speech model and inference pipeline. We also include the recently released Gemini Live API as a representative commercial model due to its user-friendly interface.
  • dGSLM [26]: A textless speech-to-speech model that generates natural conversations directly from audio. It uses (1) a HuBERT+k-means encoder [46], (2) a dual-tower Transformer for two-channel dialogue modeling, and (3) a HiFi-GAN decoder. Trained on 2,000 hours of phone calls [47], it captures both linguistic and paralinguistic cues. Originally non-interactive, we adapt it for live interaction using the official implementation [^8] https://github.com/facebookresearch/fairseq/tree/main/examples/textless_nlp/dgslm with our modifications in the Appendix.
  • Moshi [7]: A real-time speech-to-speech system combining a 7B LLM (Helium) and neural codec (Mimi) via residual vector quantization. It features an “Inner Monologue” step to improve fluency and supports overlapping speech and interruptions through a multi-stream architecture. We use the official implementation [^9] https://github.com/kyutai-labs/moshi; setup details are in the Appendix.
  • Freeze-Omni [10]: A cascaded full-duplex system with a frozen LLM pipeline. VAD triggers chunk-wise encoding, and a classification head predicts dialogue states to control turn-taking. Parallel modules handle streaming, speaking, and monitoring. Evaluated locally using the official server.
  • Gemini Live: Based on the official API documentation [^10] https://ai.google.dev/gemini-api/docs/live, we use the gemini-2.0-flash-live-001 model. The input.wav file is first converted to 16 kHz PCM-16 format, then divided into 30 ms chunks and streamed to the Gemini Live API. Server-side voice activity detection (VAD) is used to segment the audio and trigger responses. A new session is initiated after each model reply. All generated outputs are aligned with the original input duration, preserving silence in regions where no response is produced.
V Results
@@
Table III presents the results across four evaluation dimensions. The key findings are summarized as follows:
@@
Avoiding Interruptions During Speaker Pauses: All three SDMs exhibit high Takeover Rates (TOR) when managing speaker pauses, indicating frequent interruptions during natural breaks. The end-to-end models (dGSLM and Moshi) tend to interrupt more often across both real and synthetic datasets. In contrast, Freeze-Omni, which incorporates a dedicated module for predicting speaking and listening states, demonstrates a significantly lower TOR. This suggests that an explicit turn-taking control module can better manage pauses and may offer improvements if integrated into end-to-end systems. Conversely, Gemini Live achieves the lowest TOR among all models, including open-source ones, and exhibits a higher likelihood of taking over on Candor compared to synthetic data.
Backchanneling Dynamics: We evaluate using TOR, backchannel frequency (Freq), and Jensen–Shannon Divergence (JSD). Similar to pause handling, Moshi frequently takes over the turn, resulting in a high TOR. Both dGSLM and Freeze-Omni achieve lower TORs. Among the three open-sourced SDMs, dGSLM produces the most backchannel responses with more natural timing (Freq = 0.015, JSD = 0.934) when it does not take over the turn. In contrast, Freeze-Omni remains largely silent, producing few backchannel responses. The commercial Gemini Live achieves the lowest TOR and the best JSD, indicating superior ability in identifying appropriate moments for backchanneling.
Latency of Turn-Taking: We assessed TOR and average response latency on the Candor dataset. dGSLM and Moshi exhibit high TORs and respond quickly, with an average latency of around 0.3 seconds. Freeze-Omni, due to its cascaded architecture—which first generates text and then synthesizes speech—exhibits higher latency. Its lower TOR likely reflects missed opportunities to take over the turn, possibly due to failures in detecting turn ends. Interestingly, Gemini Live achieves a TOR of only 0.655, suggesting that even commercial models sometimes fail to take the turn in real dialogue data.
@@
Managing User Interruptions: Freeze-Omni handles user interruptions effectively, achieving significantly higher contextual relevance scores while maintaining acceptable latency. This strength is attributed to its ’model-as-a-server’ strategy, which leverages a pool of models to manage user barge-ins efficiently. Gemini Live shows comparable performance, with relatively better TOR and latency, though it yields a slightly lower GPT-4o score. In contrast, the end-to-end SDMs struggle with coherence. Moshi responds promptly but yields a lower GPT-4o score (0.765), while dGSLM performs poorly, with high latency and diminished content quality (GPT-4o score: 0.201). These results highlight the challenges end-to-end systems face in preserving semantic coherence during user interruptions.
VI Conclusion
In this paper, we introduced Full-Duplex-Bench, a benchmark designed to evaluate critical aspects of full-duplex spoken dialogue models. Our framework targets key interaction dimensions—pause handling, backchanneling, smooth turn-taking, and user interruption management—addressing limitations of existing benchmarks that primarily focus on half-duplex settings or coarse corpus-level metrics. To ensure systematic and reproducible evaluation, we propose automatic metrics tailored to real-time interaction. Experiments on open-source full-duplex models reveal distinct model features and highlight areas for improvement. We hope Full-Duplex-Bench provides a practical foundation for evaluating full-duplex spoken dialogue systems. By releasing our data and behavior-specific metrics, we aim to enable fair comparison and advance research toward more natural and responsive human-computer interaction.